# Natal kick by early-asymmetrical pairs of jets to the neutron star of supernova remnant S147

### v1

@@ 4 Summary
We studied the CCSNR S147, which exhibits a bipolar morphology characterized by a prominent pair of large ears, and related its NS velocity to its morphology. Based on the wide morphology of the ears rather than the more common conical structure and their internal structures, in Section 2.1 we proposed that the ears were inflated by two pairs of jets (or a precessing pair of jets). We draw the two axes of the two pairs of jets on the upper panel of Figure 1 by dashed yellow lines. This point-symmetric morphology is completely compatible with the JJEM and challenges the neutrino-driven explosion mechanism. To connect the NS kick velocity to the morphology of S147, we define one axis through the ears and between the two jet axes; this is the main jet axis of S147 (solid yellow line in Figures 1 and 2). The projected intersection between the ears symmetry axis and the NS trajectory allowed us to estimate the age of the NS-SNR system to be (Section 2.1 and Appendix A).
@@
In the kick-BEAP mechanism, the kick is along the axis of the pair of jets, which is the momentary angular momentum of the accreted mass onto the NS. The pair of jets might compress two opposite zones, as we suggest is the case with the two zones on the ends of the cyan lines on Figure 2, or inflate ears (bubbles), as the two ears at the ends of the yellow line on Figure 2. We attribute the shaping of the pair of ears to two pairs of jets, two of several that exploded the star in the frame of the JJEM. We note that, in principle, the ears could have been formed by pre-explosion mass loss, and the ejecta could have expanded into the CSM that already had two ears, as suggested for some type Ia SNRs (e.g., Tsebrenko & Soker (2015); Soker (2022c, 2024h)). We note that in the two Ia SNRs with prominent ears, Kepler SNR and SNR G1.9+0.3, the ears are approximately equal in size and do not exhibit any signatures within the main SNR. In most CCSNRs with pairs of large ears, the two opposite ears are not equal, and in some cases, they have signatures that extend deep into the main shell, e.g., SNR G309.2-00.6 (for an image, e.g., Gaensler et al. 1998) and the Vela SNR (for discussion see Soker & Shishkin 2025, and for an image Mayer et al. 2023). In most CCSNe, the axis of the CSM is not aligned with the axis of the ejecta (e.g., discussion by Vasylyev et al. (2025)), as is most clearly evident in SN 1987A.
@@
Acknowledgements
We thank Mr. Christian Koll for allowing us to use his image of S147. N.S. thanks the Charles Wolfson Academic Chair at the Technion for the support.
The eROSITA data shown here were processed using the eSASS software system developed by the German eROSITA consortium.
@@ Acknowledgements
Funding Statement
A grant from the Asher Space Research Institute at the Technion and a grant from the Pazy Foundation supported this research.
Appendix A Appendix A: Age estimate
@@
Upper panel: Combined X-ray and image of S147 (as Fig 2d). As in Figures 1,2, we mark the main symmetry axis connecting the eastern and western ears with a solid-yellow line, and the two sub-ears with dashed-yellow lines. We outline the approximate extents of the two sub-ear structures with solid-gray lines. Black dots (composing a black tetragon) denote intersection points between lines connecting the sub-ears, originating along the extend of the smaller ears, and blue dots denote the intersection points that fall in the vicinity (within ) of the NS trajectory (on the plain of the sky), marked with a dashed-red line.
Lower panel: A histogram of the intersection points from the upper panel (light blue), along the NS trajectory in , as derived by the proper motion estimate of Yao et al. (2021). We take the median and 15th and 85th percentiles as our estimate and confidence intervals, respectively. We mark the other morphological age estimates of Yao et al. (2021) in red, and of Kramer et al. (2003) in orange. Other age estimates from the remnant expansion rate (Kundu et al. 1980, Sofue et al. 1980) and NS spin-down rate (Anderson et al. 1996) are longer, and are marked as arrows on the left.
We draw the histogram of these triple-intersecting points, of the two symmetry axes, and the proper motion direction, in the lower panel of Figure A.1. The median is selected as the proposed age (solid-green line), and the 15th and 85th percentiles (corresponding to the edges of the Tophat-like distribution) as our confidence interval (dashed-green lines). The vertical axis is an arbitrary number of points count (scales with number of initial points considered, increased to get a sufficient statistic; ‘Freedman–Diaconis’ binned). The horizontal axis is transformed from pixel distance of the points from the NS current location along the trajectory to a proposed age using the proper motion velocity. We find the age to be , and denote this age along the NS trajectory in the upper panel with a green line and dot. In the lower panel, we also display other age estimates for the explosion; these vary in method, but can be categorized into three classes:
NS Spin-down:
Anderson et al. (1996):

### v2

@@ 4 Summary
We studied the CCSNR S147, which exhibits a bipolar morphology characterized by a prominent pair of large ears, and related its NS velocity to its morphology. Based on the wide morphology of the ears rather than the more common conical structure and their internal structures, in Section 2.1 we proposed that the ears were inflated by two pairs of jets (or a precessing pair of jets). We draw the two axes of the two pairs of jets on the upper panel of Figure 1 by dashed yellow lines. This point-symmetric morphology is completely compatible with the JJEM and challenges the neutrino-driven explosion mechanism.
To connect the NS kick velocity to the morphology of S147, we define one axis through the ears and between the two jet axes; this is the main jet axis of S147 (solid yellow line in Figures 1 and 2). The projected intersection between the ears symmetry axis and the NS trajectory allowed us to estimate the age of the NS-SNR system to be (Section 2.1 and Appendix A).
@@
In the kick-BEAP mechanism, the kick is along the axis of the pair of jets, which is the momentary angular momentum of the accreted mass onto the NS. The pair of jets might compress two opposite zones, as we suggest is the case with the two zones on the ends of the cyan lines on Figure 2, or inflate ears (bubbles), as the two ears at the ends of the yellow line on Figure 2. We attribute the shaping of the pair of ears to two pairs of jets, two of several that exploded the star in the frame of the JJEM. We note that, in principle, the ears could have been formed by pre-explosion mass loss, and the ejecta could have expanded into the CSM that already had two ears, as suggested for some type Ia SNRs (e.g., Tsebrenko & Soker 2015; Soker 2022c, 2024g). We note that in the two Ia SNRs with prominent ears, Kepler SNR and SNR G1.9+0.3, the ears are approximately equal in size and do not exhibit any signatures within the main SNR. In most CCSNRs with pairs of large ears, the two opposite ears are not equal, and in some cases, they have signatures that extend deep into the main shell, e.g., SNR G309.2-00.6 (for an image, e.g., Gaensler et al. 1998) and the Vela SNR (for discussion see Soker & Shishkin 2025, and for an image Mayer et al. 2023). In most CCSNe, the axis of the CSM is not aligned with the axis of the ejecta (e.g., discussion by Vasylyev et al. 2025), as is most clearly evident in SN 1987A.
@@
Appendix A Appendix: Age estimate
@@
We draw the histogram of these triple-intersecting points, of the two symmetry axes, and the proper motion direction, in the lower panel of Figure A.1. The median is selected as the proposed age (solid-green line), and the 15th and 85th percentiles (corresponding to the edges of the Tophat-like distribution) as our confidence interval (dashed-green lines). The vertical axis is the intersection points density (‘Freedman–Diaconis’ binned). The horizontal axis is transformed from pixel distance of the points from the NS current location along the trajectory to a proposed age using the proper motion velocity. We find the age to be , and denote this age along the NS trajectory in the upper panel with a green line and dot. In the lower panel, we also display other age estimates for the explosion; these vary in method, but can be categorized into three classes: